OpenStudy (anonymous):

A triangle has coordinates A (1, 5), B (-2, 1) and C (0, -4). What are the new coordinates if the triangle is dilated with a scale factor of 1/5 ?

OpenStudy (johnweldon1993):

So it's as easy as taking each coordinate and multiplying them by 1/5

OpenStudy (anonymous):

each number in the coordinate by 1/5 ?? @johnweldon1993

OpenStudy (johnweldon1993):

Right, for example \[\large A = (1,5)\] would be \[\large (1 \times \frac{1}{5} , 5 \times \frac{1}{5})\] or \[\large (\frac{1}{5} , 1)\]

OpenStudy (anonymous):

thank you so much !!

OpenStudy (johnweldon1993):

Of course :)